# Robust Radiotherapy Planning with Spatially Based Uncertainty Sets

### v1

@@ Abstract
Keywords: Radiotherapy planning, robust optimization, biomarker uncertainty, row and column generation
@@ 1 Introduction
Radiotherapy treatment, and in particular intensity modulated radiation therapy (IMRT), involves setting up an array of beams with variable intensities to irradiate a tumor. Planning such a treatment entails selecting the intensity of each beam or beam unit (called a beamlet). To facilitate such planning, the effect of the beam intensities – that is, the resulting radiation dose (whose unit of measure is 1 Gy = 1 J/kg) – on both the tumor and the healthy surrounding organs must be analyzed. The dose is measured over a three-dimensional grid of spatial cube units called voxels. The size of a voxel, and the granularity of the partition into voxels, are determined by the radiation beamlet size and the imaging resolution. Recent technology allows for voxel edge lengths of 2-3 mm. The dose at each voxel is generally assumed to be linearly related to the chosen beamlet intensities. In general, this is the theme of radiotherapy treatment planning (RTP), which may prescribe the treatment over both time and space. The particular optimization problem for setting up the beamlet intensities is known as fluence map optimization (FMO). A variety of formulations have been proposed for this problem. Typically, such formulations require that the tumor voxels are sufficiently irradiated, while healthy tissue voxels are spared from receiving an unhealthy dose. The formulations differ in the way that these requirements are handled. Additional clinical requirements can also be factored into the model. Some formulations aim to maximize the average tumor voxel dose (Preciado-Walters et al. 2006), while constraining the dose received by healthy organs. Other formulations minimize a measure of the dose received by healthy tissue voxels, or by all voxels, subject to ensuring that target (i.e., tumor) voxels receive at least the prescribed dose (Romeijn et al. 2006).
@@
Our approach is to maximize the tumor’s “weakest link”: the voxels receiving the minimum dose of all tumor voxels. This is a widely accepted objective for evaluating radiotherapy planning (Deasy 1997, Yan et al. 1997, ICR 2010, Preciado-Walters et al. 2004), subject to constraints on the healthy organ dose and on tumor dose homogeneity. In particular, the objective of maximizing the minimum dose objective is argued to better support biological optimization, as “a conservative surrogate for biological effect”, compared with minimization of deviations from prescribed doses (Deasy 1997). The importance of combining this objective with homogeneity constraints is outlined by Yan et al. (1997), Allen Li et al. (2012), Khairi et al. (2021).
@@
Advances in imaging have exposed uncertainties in the effect of the prescribed radiation dosage. It is now known that tumors are not homogeneous, but rather, the sensitivity of cells varies across the tumor and changes over time (Bentzen and Gregoire 2011, Saka et al. 2014, Titz and Jeraj 2008, Nohadani and Roy 2017). In an effort to improve patient outcomes, there is a need for the model to account for the uncertain problem data. Modeling uncertainty in an optimization problem usually results in added complexity and scale. Uncertainty is involved in all radiotherapy treatment planning stages, starting with the target volume definition, but also in relation to the dose required to control the tumor, healthy tissue tolerances, and the actual dose delivered to the patient. Uncertainty in radiotherapy planning has been extensively addressed in the literature, mainly in the context of geometric uncertainty that results from the patient’s movement and from imaging inaccuracies; see, for example, Bortfeld et al. (2008), the recent survey by Unkelbach et al. (2018), and the references therein.
@@
Approaches for dealing with uncertainty in FMO include stochastic programming, robust optimization and distributionally robust optimization. The stochastic approach assumes that the distribution of the uncertainty is known and its parameters can be reasonably estimated. In practice, estimating the exact distribution is difficult, and incorrectly estimating the distribution may lead to solutions that are not clinically viable. In contrast, the robust optimization methodology (Ben-Tal et al. 2009) does not assume any knowledge of the distribution, but rather, defines a set that contains all possible uncertainty realizations that need to be hedged against. By making few or no assumptions about the distribution, the robust approach optimizes the treatment for the worst-case realization of the uncertainty. Motivated by motion uncertainty, Chu et al. (2006) propose a robust formulation with uncertain influence matrices. The distributionally robust approach combines the stochastic and robust approaches – it protects against the worst-case distribution within a family of possible probability distributions of the uncertain parameters. For example, Bortfeld et al. (2008) model motion uncertainty by defining families of distributions that correspond to possible motion patterns of patients. Unkelbach et al. (2007) consider a (discrete) uncertainty set for beamlet range uncertainties in proton therapy (IMPT). This uncertainty set is constructed so that beamlets of the same beam must reach the same range. In this work they consider the objective of minimizing a weighted sum of the deviations from a prescribed dose. However, the model being solved may be overly conservative as the worst case is applied to each of the voxels separately, rather than to the objective. A similar scenario-based approach to account for range uncertainty and setup uncertainty in IMPT is also used in Pflugfelder et al. (2008).
@@
Another source of uncertainty, also mentioned in Unkelbach et al. (2018), arises from the biological conditions of the patients, which the authors consider a new frontier of robust RTP optimization. These conditions may affect the radiosensitivity of the tumor to the dose. The resulting effective dose is referred to as the biologically-adjusted dose. We note that this adjustment is different from the biological effective dose (BED) often modeled by the linear-quadratic (LQ) model (Jones et al. 2001), which aims to estimate the biological effectiveness in terms of cell-kill probability for a given physical dose. In contrast, the biologically adjusted dose corrects the physical dose to account for the impact of a biological condition on the effective dose at a particular voxel. For the case of oxygen levels (or the lack thereof also known as hypoxia), this is described by the oxygen enhancement ratio (OER), which is the ratio between doses required to obtain the same biological effect (Hill et al. 2015, McMahon 2018), or its normalized variant known as the oxygen modification factor (OMF) (Titz and Jeraj 2008). Information about a living tissue’s (e.g., a tumor’s) biological conditions is available through bio-markers and imaging data. Specifically, the tumor’s oxygenation level, which strongly affects its radiosensitivity, is measured by imaging techniques such as fluoromisonidazole (FMISO) (Toma-Dasu et al. 2012). Translating the imaging data into radiosensitivity values is a crude estimation that is subject to uncertainty. Further, the oxygenation level, and accordingly the radiosensitivity, may change between the time of the image scan and the commencement of treatment. In fact, biological uncertainty in radiation therapy, and the effect of oxygenation on radiosensitivity in particular, is an active area of research in the medical physics and other research communities (Titz and Jeraj 2008, Toma-Dasu et al. 2012, Shang et al. 2021). Biological conditions and radiosensitivities are considered in the context of FMO by Saka et al. (2014) but their proposed methods do not attempt to model and protect from the associated uncertainty. Robust static (e.g., cumulative) dose planning under biological uncertainty has been proposed for biological uncertainty by Li et al. (2015) and Ten Eikelder et al. (2020). Ten Eikelder et al. (2020) propose a tumor control probability (TCP) based model assuming voxel independence and simple box uncertainty for its parameters. Simple box uncertainty here implies conservative solutions where all voxels are assumed to be simultaneously insensitive. Li et al. (2015) propose a dose-based model with budgeted uncertainty (see Bertsimas and Sim (2004)) on the deviation of the voxels’ prescribed doses from their nominal value. In this model, the conservatism is controlled by a single budget parameter that indicates how many voxel doses can deviate. However, in this model the budget is allocated linearly to the change in the prescribed doses, so it might not accurately capture measurement errors whose effect on radiosensitivity may be highly nonlinear Titz and Jeraj (2008). Also, this model does not limit the differences in radiosensitivities among neighboring voxels.
@@
Recent papers that consider radiosensitivity uncertainty in adaptive radiation therapy planning include Nohadani and Roy (2017), Roy et al. (2022), Jeyakumar et al. (2023). These papers assume a perscribed dose approach where the objective is to minimize the total dose administered while satisfying simple lower and upper dose bound constraints on target voxels and OARs, respectively. Consequently, the biological uncertainty in this formulation appears only on the right-hand side. Nohadani and Roy (2017) utilize simple box uncertainty sets containing the measured radiosensitivities, and model the evolution of this uncertainty through time. In the current paper, we also formulate and solve a robust FMO that accounts for biological uncertainty at the voxel level where the main focus is to model the spatial connectivity or dependence of the voxel uncertainties. Both Roy et al. (2022), Jeyakumar et al. (2023) propose models similar to Nohadani and Roy (2017). Roy et al. (2022) studies an extension of the model to a multistage adaptive problem, while Jeyakumar et al. (2023) consider adaptivity to uncertainty in influence matrix resulting in a semidefinite program.
@@
In contrast to previous work, our approach is to model and explicitly account for the spatial interdependence among individual voxel radiosenstivity uncertainties. Accounting for such interdependence is expected to improve over simpler robust approaches that tend to result in overly conservative solutions. Spatial dependence has been considered in other applications; for example by estimating covariance matrices and applying these using ellipsoid uncertainty sets in transportation Chassein et al. (2019), or general polyhedral uncertainty sets in energy Lorca and Sun (2014). These previous approaches may be less appropriate for biological uncertainty in radiotherapy because there may not be sufficient data to reliability estimate the correlations between voxel radiosensitivities. Also, applying both types of uncertainty sets to large-scale radiotherapy optimization problems, may be computationally intractable in practice. In radiotherapy, the scenario-based approaches to range and setup uncertainties in IMPT appear to resemble robust optimization with spatial uncertainty (for example Unkelbach et al. (2007), Pflugfelder et al. (2008) discussed above). However, the uncertainty applies to the beamlets rather than to voxels and it is addressed by scenario-based approaches, so the dependency is not explicitly modeled. In contrast, the current paper focuses on spatial uncertainty modelling, reformulation of the corresponding robust FMO problem and solution methods for solving this problem. Dividing the administered dose into fractions and accounting for treatment sessions over time (also known as fractions) may involve a more elaborate uncertainty model over both space and time that is left for future work.
@@
To make our proposed model applicable to clinical settings, we also consider important dose-volume constraints. These constraints limit the percentage of healthy-organ voxels that can receive more than a specified radiation dose. These are hard combinatorial constraints, which are virtually impossible to handle exactly given the problem size. Hence, many solution approaches have been proposed to approximately solve such constraints in the literature (Ferris et al. 2003, Preciado-Walters et al. 2006, Romeijn et al. 2006). However, effectively solving models that incorporate such constraints remains a challenging problem, especially in the context of the large scale problems considered herein.
@@ 2.1 Model Introduction - Nominal Model
We also consider homogeneity constraints that limit the dose range across the PTV (Preciado-Walters et al. 2004). This is intended to prevent the creation of “hot spots” in the tumor tissue, which cause damage to the healthy tissue in which the tumor is embedded. Specifically, our homogeneity constraints are of the form $\mu f({\bf d})-\phi_{v}{d}_{v}\geq 0$, for all $v\in T$, where $\mu>1$ is a given homogeneity parameter. Homogeneity has been traditionally applied to the physical dose Allen Li et al. (2012), Kataria et al. (2012), Khairi et al. (2021). The possibility of applying it to the adjusted dose has been enabled by recent technological advances in estimating necessary biological information, specifically the radiosensitivity vector $\phi$. The main goal of the current paper is to address the challenge of modeling the uncertainty in $\phi$ and using this information to improve radiotherapy treatment planning. The optimization problem of determining the intensity of the beams given the radiosensitivity parameter $\boldsymbol{\phi}$ is formulated as
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (1b)
subject to $\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (1c)
$\mu\underaccent{\bar}{d}-\phi_{v}d_{v}(x)\geq 0,$ $v\in T$ (1d)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k},$ (1e)
@@ 2.2 Robust Optimization Model with Spatially Bound Uncertainty Set
We now present a modeling scheme to deal with the uncertainty in the parameter vector $\boldsymbol{\phi}$. The radiosensitivity uncertainty is caused by multiple factors, including imaging noise, biomarker conversion error, and variation over time. Inferring a distribution of the radiosensitivity parameter based on limited patient-specific (personalized) data is nontrivial and prone to error. Moreover, maximizing the clinical outcome (in our model, maximizing the minimum PTV adjusted dose) while using a high-dimensional multivariate probability distribution to model the biological uncertainty, would result in chance-constrained optimization problem (Luedtke et al. 2010), which is notorious for being computationally hard. Instead, our approach to modeling uncertainty follows the robust optimization (RO) methodology. In this modeling framework, the uncertain parameters are assumed to lie in a predetermined uncertainty set denoted by $\mathcal{U}$. An RO problem formulation is designed to protect its solutions against all realizations in the uncertainty set. In particular, an optimal solution $({\bf x},\underaccent{\bar}{d})$ must remain feasible for every possible realization $\boldsymbol{\phi}\in\mathcal{U}$ while maximizing the worst-case (minimum) PTV adjusted dose.
Formulating the robust counterpart (RC) of (2.1), in which variable $\underline{d}$ now depends on the uncertain $\phi$ subject to uncertain homogeneity constraints (1d), would lead to an adjustable RO problem. Such problems are computationally intensive (Ben-Tal et al. 2004), and although it may be possible to solve these adjustable formulations when the problem is moderately sized, we find it more effective to avoid the adjustable robust model by reformulating the homogeneity constraints according to the equivalence
@@
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (3b)
subject to $\min_{\boldsymbol{\phi}\in\mathcal{U}}\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (3c)
$\max_{\boldsymbol{\phi}\in\mathcal{U}}\left\{\phi_{v}d_{v}(x)-\mu\phi_{u}d_{u}(x)\right\}\leq 0,$ $u,v\in T$ (3d)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (3e)
@@
Given a radiosensitivity vector $\widehat{\boldsymbol{\phi}}$ that is estimated based on a patient’s imaging data at the start of the treatment, the goal is to construct an uncertainty set that satisfies our three modeling requirements, described in Section 2.2. Perhaps the simplest uncertainty set that can be applied is a box (in particular, a hypercube) uncertainty set, which allows the radiosensitivity of each PTV voxel $v$ to deviate from its calculated value $\widehat{\phi}_{v}$ by at most $\delta>0$, where $\delta$ is a given parameter. That is, $\mathcal{U}_{\text{B}}=\{\boldsymbol{\phi}\in\mathbb{R}^{|T|}:|\phi_{v}-\widehat{\phi}_{v}|\leq\delta,\;\forall v\in T\}$. This uncertainty set, used for example in Nohadani and Roy (2017), admits a straightforward robust formulation. However, if $\delta$ is chosen to be sufficiently large to capture the actual uncertainty in the tumor biological conditions, then the resulting set may be overly conservative, thereby violating one of our key modeling requirements (see Section 2.2). To restrict the uncertainty further, a key observation to be made is that voxels that are situated close to each other should have similar radiosensitivities. In Figure 1, both plots demonstrate that the radiosensitivity differences are small for neighboring voxels and increase with the distance.
@@
Figure 1: Relationship between radiosensitivity difference and pairwise PTV voxel distance in the first-visit data of Patient 1 from the TCIA brain dataset (Clark et al. 2013).
Incorporating this observation into the uncertainty model leads to a spatially bound (SB) uncertainty set; in particular, we define
@@
For any $u,v\in T$, $\gamma_{uv}=\Gamma(\operatorname{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dist}}(u,v))$, where $\operatorname{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dist}}(u,v)$ is the distance in voxels between $u$ and $v$, and $\Gamma:\mathbb{R}\rightarrow[0,1]$ is a nondecreasing subadditive function satisfying $\Gamma(x)=0$ if and only if $x=0$.
@@
The $\gamma_{uv}$ parameters for $u,v\in T$ can be personalized to a particular patient’s data, thereby addressing requirement (i). Requirement (iii) may be satisfied by imposing the spatial constraints associated with the parameters $\gamma_{uv}$, for $u,v\in T$, to augment the box constraints in the definition of the spatially bound uncertainty set. In the absence of these spatial constraints, the uncertainty set would reduce to the overly conservative box uncertainty set. The following section addresses the question of whether optimal solutions of (2.2) for uncertainty set $\mathcal{U}_{\text{SB}}$ can be computed efficiently in practice (requirement (ii)).

### v2

@@ Abstract
Keywords: Radiotherapy planning; robust optimization; biomarker uncertainty; row and column generation.
@@ 1 Introduction
Radiotherapy treatment, and in particular intensity modulated radiation therapy (IMRT), involves setting up an array of beams with variable intensities to irradiate a tumor. Planning such a treatment entails selecting the intensity of each beam or beam unit (called a beamlet). To facilitate such planning, the effect of the beam intensities – that is, the resulting radiation dose (whose unit of measure is 1 Gy = 1 J/kg) – on both the tumor and the healthy surrounding organs must be analyzed. The dose is measured over a three-dimensional grid of spatial cube units called voxels. The size of a voxel, and the granularity of the partition into voxels, are determined by the radiation beamlet size and the imaging resolution. Recent technology allows for voxel edge lengths of 2-3 mm. The dose at each voxel is generally assumed to be linearly related to the chosen beamlet intensities. In general, this is the theme of radiotherapy treatment planning (RTP), which may prescribe the treatment over both time and space. The particular optimization problem for setting up the beamlet intensities is known as fluence map optimization (FMO). A variety of formulations have been proposed for this problem. Typically, such formulations require that the tumor voxels are sufficiently irradiated, while healthy tissue voxels are spared from receiving an unhealthy dose. The formulations differ in the way that these requirements are handled. Additional clinical requirements can also be factored into the model. Some formulations aim to maximize the average tumor voxel dose (Preciado-Walters et al., 2006), while constraining the dose received by healthy organs. Other formulations minimize a measure of the dose received by healthy tissue voxels, or by all voxels, subject to ensuring that target (i.e., tumor) voxels receive at least the prescribed dose (Romeijn et al., 2006).
@@
Our approach is to maximize the tumor’s “weakest link”: the voxels receiving the minimum dose of all tumor voxels. This is a widely accepted objective for evaluating radiotherapy planning (Deasy, 1997, Yan et al., 1997, ICR, 2010, and Preciado-Walters et al., 2004), subject to constraints on the healthy organ dose and on tumor dose homogeneity. In particular, the objective of maximizing the minimum dose objective is argued to better support biological optimization, as “a conservative surrogate for biological effect”, compared with minimization of deviations from prescribed doses (Deasy, 1997). The importance of combining this objective with homogeneity constraints is outlined by Yan et al., (1997), Allen Li et al., (2012), and Khairi et al., (2021).
@@
Advances in imaging have exposed uncertainties in the effect of the prescribed radiation dosage. It is now known that tumors are not homogeneous, but rather, the sensitivity of cells varies across the tumor and changes over time (Bentzen and Gregoire, 2011, Saka et al., 2014, Titz and Jeraj, 2008, and Nohadani and Roy, 2017). In an effort to improve patient outcomes, there is a need for the model to account for the uncertain problem data. Modeling uncertainty in an optimization problem usually results in added complexity and scale. Uncertainty is involved in all radiotherapy treatment planning stages, starting with the target volume definition, but also in relation to the dose required to control the tumor, healthy tissue tolerances, and the actual dose delivered to the patient. Uncertainty in radiotherapy planning has been extensively addressed in the literature, mainly in the context of geometric uncertainty that results from the patient’s movement and from imaging inaccuracies; see, for example, Bortfeld et al., (2008), the recent survey by Unkelbach et al., (2018), and the references therein.
@@
Approaches for dealing with uncertainty in FMO include stochastic programming, robust optimization and distributionally robust optimization. The stochastic approach assumes that the distribution of the uncertainty is known and its parameters can be reasonably estimated. In practice, estimating the exact distribution is difficult, and incorrectly estimating the distribution may lead to solutions that are not clinically viable. In contrast, the robust optimization methodology (Ben-Tal et al., 2009) does not assume any knowledge of the distribution, but rather, defines a set that contains all possible uncertainty realizations that need to be hedged against. By making few or no assumptions about the distribution, the robust approach optimizes the treatment for the worst-case realization of the uncertainty. Motivated by motion uncertainty, Chu et al., (2006) propose a robust formulation with uncertain influence matrices. The distributionally robust approach combines the stochastic and robust approaches – it protects against the worst-case distribution within a family of possible probability distributions of the uncertain parameters. For example, Bortfeld et al., (2008) model motion uncertainty by defining families of distributions that correspond to possible motion patterns of patients. Unkelbach et al., (2007) consider a (discrete) uncertainty set for beamlet range uncertainties in proton therapy (IMPT). This uncertainty set is constructed so that beamlets of the same beam must reach the same range. In this work they consider the objective of minimizing a weighted sum of the deviations from a prescribed dose. However, the model being solved may be overly conservative as the worst case is applied to each of the voxels separately, rather than to the objective. A similar scenario-based approach to account for range uncertainty and setup uncertainty in IMPT is also used in Pflugfelder et al., (2008).
@@
Another source of uncertainty, also mentioned in Unkelbach et al., (2018), arises from the biological conditions of the patients, which the authors consider a new frontier of robust RTP optimization. These conditions may affect the radiosensitivity of the tumor to the dose. The resulting effective dose is referred to as the biologically-adjusted dose. We note that this adjustment is different from the biological effective dose (BED) often modeled by the linear-quadratic (LQ) model (Jones et al., 2001), which aims to estimate the biological effectiveness in terms of cell-kill probability for a given physical dose. In contrast, the biologically adjusted dose corrects the physical dose to account for the impact of a biological condition on the effective dose at a particular voxel. For the case of oxygen levels (or the lack thereof also known as hypoxia), this is described by the oxygen enhancement ratio (OER), which is the ratio between doses required to obtain the same biological effect (Hill et al., 2015 and McMahon, 2018), or its normalized variant known as the oxygen modification factor (OMF) (Titz and Jeraj, 2008). Information about a living tissue’s (e.g., a tumor’s) biological conditions is available through bio-markers and imaging data. Specifically, the tumor’s oxygenation level, which strongly affects its radiosensitivity, is measured by imaging techniques such as fluoromisonidazole (FMISO) (Toma-Dasu et al., 2012). Translating the imaging data into radiosensitivity values is a crude estimation that is subject to uncertainty. Further, the oxygenation level, and accordingly the radiosensitivity, may change between the time of the image scan and the commencement of treatment. In fact, biological uncertainty in radiation therapy, and the effect of oxygenation on radiosensitivity in particular, is an active area of research in the medical physics and other research communities (Titz and Jeraj, 2008, Toma-Dasu et al., 2012, and Shang et al., 2021). Biological conditions and radiosensitivities are considered in the context of FMO by Saka et al., (2014) but their proposed methods do not attempt to model and protect from the associated uncertainty. Robust static (e.g., cumulative) dose planning under biological uncertainty has been proposed by Li et al., (2015) and Ten Eikelder et al., (2020). Ten Eikelder et al., (2020) propose a tumor control probability (TCP) based model assuming voxel independence and simple box uncertainty for its parameters. Simple box uncertainty here implies conservative solutions where all voxels are assumed to be simultaneously insensitive. Li et al., (2015) propose a dose-based model with budgeted uncertainty (see Bertsimas and Sim, (2004)) on the deviation of the voxels’ prescribed doses from their nominal value. In this model, the conservatism is controlled by a single budget parameter that controls how many voxels have prescribed doses that (maximally) deviate from their nominal values.. However, in this model the budget is allocated linearly to the change in the prescribed doses, so it might not accurately capture measurement errors whose effect on radiosensitivity may be highly nonlinear (Titz and Jeraj, 2008). Also, this model does not limit the differences in radiosensitivities among neighboring voxels, and thus may consider unrealistic scenarios where the dose in one voxel radiosensitivity is significantly lower than its neighboring voxels.
@@
Recent papers that consider radiosensitivity uncertainty in adaptive radiation therapy planning include Nohadani and Roy, (2017), Roy et al., (2022), and Jeyakumar et al., (2023). These papers assume a perscribed dose approach where the objective is to minimize the total dose administered while satisfying simple lower and upper dose bound constraints on target voxels and OARs, respectively. Consequently, the biological uncertainty in this formulation appears only on the right-hand side. Nohadani and Roy, (2017) utilize simple box uncertainty sets containing the measured radiosensitivities, and model the evolution of this uncertainty through time. In the current paper, we also formulate and solve a robust FMO that accounts for biological uncertainty at the voxel level where the main focus is to model the spatial connectivity or dependence of the voxel uncertainties. Both Roy et al., (2022) and Jeyakumar et al., (2023) propose models similar to Nohadani and Roy, (2017). Roy et al., (2022) studies an extension of the model to a multistage adaptive problem, while Jeyakumar et al., (2023) consider adaptivity to uncertainty in influence matrix resulting in a semidefinite program.
@@
In contrast to previous work, our approach is to model and explicitly account for the spatial interdependence among individual voxel radiosenstivity uncertainties. Accounting for such interdependence is expected to improve over simpler robust approaches that tend to result in overly conservative solutions. Spatial dependence has been considered in other applications; for example by estimating covariance matrices and applying these using ellipsoid uncertainty sets in transportation (Chassein et al., 2019), or general polyhedral uncertainty sets in energy (Lorca and Sun, 2014). These previous approaches may be less appropriate for biological uncertainty in radiotherapy because there may not be sufficient data to reliability estimate the correlations between voxel radiosensitivities. Also, applying both types of uncertainty sets to large-scale radiotherapy optimization problems, may be computationally intractable in practice. In radiotherapy, the scenario-based approaches to range and setup uncertainties in IMPT appear to resemble robust optimization with spatial uncertainty (for example Unkelbach et al., (2007) and Pflugfelder et al., (2008) discussed above). However, the uncertainty applies to the beamlets rather than to voxels and it is addressed by scenario-based approaches, so the dependency is not explicitly modeled. In contrast, the current paper focuses on spatial uncertainty modelling, reformulation of the corresponding robust FMO problem and solution methods for solving this problem. Dividing the administered dose into fractions and accounting for treatment sessions over time (also known as fractions) may involve a more elaborate uncertainty model over both space and time that is left for future work.
@@
To make our proposed model applicable to clinical settings, we also consider important dose-volume constraints. These constraints limit the percentage of healthy-organ voxels that can receive more than a specified radiation dose. These are hard combinatorial constraints, which are virtually impossible to handle exactly given the problem size. Hence, many solution approaches have been proposed to approximately solve such constraints in the literature (Ferris et al., 2003, Preciado-Walters et al., 2006, and Romeijn et al., 2006). However, effectively solving models that incorporate such constraints remains a challenging problem, especially in the context of the large scale problems considered herein.
@@ 2.1 Model Introduction - Nominal Model
We also consider homogeneity constraints that limit the dose range across the PTV (Preciado-Walters et al., 2004). This is intended to prevent the creation of “hot spots” in the tumor tissue, which cause damage to the healthy tissue in which the tumor is embedded. Specifically, our homogeneity constraints are of the form $\mu f({\bf d})-\phi_{v}{d}_{v}\geq 0$, for all $v\in T$, where $\mu>1$ is a given homogeneity parameter. Homogeneity has been traditionally applied to the physical dose (Allen Li et al., 2012, Kataria et al., 2012, and Khairi et al., 2021). The possibility of applying it to the adjusted dose has been enabled by recent technological advances in estimating necessary biological information, specifically the radiosensitivity vector $\phi$. The main goal of the current paper is to address the challenge of modeling the uncertainty in $\phi$ and using this information to improve radiotherapy treatment planning. The optimization problem of determining the intensity of the beams given the radiosensitivity parameter $\boldsymbol{\phi}$ is formulated as
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (1a)
subject to $\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (1b)
$\mu\underaccent{\bar}{d}-\phi_{v}d_{v}(x)\geq 0,$ $v\in T$ (1c)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k},$ (1d)
@@ 2.2 Robust Optimization Model with Spatially Bound Uncertainty Set
We now present a modeling scheme to deal with the uncertainty in the parameter vector $\boldsymbol{\phi}$. The radiosensitivity uncertainty is caused by multiple factors, including imaging noise, biomarker conversion error, and variation over time. Inferring a distribution of the radiosensitivity parameter based on limited patient-specific (personalized) data is nontrivial and prone to error. Moreover, maximizing the clinical outcome (in our model, maximizing the minimum PTV adjusted dose) while using a high-dimensional multivariate probability distribution to model the biological uncertainty, would result in chance-constrained optimization problem (Luedtke et al., 2010), which is notorious for being computationally hard. Instead, our approach to modeling uncertainty follows the robust optimization (RO) methodology. In this modeling framework, the uncertain parameters are assumed to lie in a predetermined uncertainty set denoted by $\mathcal{U}$. An RO problem formulation is designed to protect its solutions against all realizations in the uncertainty set. In particular, an optimal solution $({\bf x},\underaccent{\bar}{d})$ must remain feasible for every possible realization $\boldsymbol{\phi}\in\mathcal{U}$ while maximizing the worst-case (minimum) PTV adjusted dose.
Formulating the robust counterpart (RC) of (1), in which variable $\underline{d}$ now depends on the uncertain $\phi$ subject to uncertain homogeneity constraints (1c), would lead to an adjustable RO problem. Such problems are computationally intensive (Ben-Tal et al., 2004), and although it may be possible to solve these adjustable formulations when the problem is moderately sized, we find it more effective to avoid the adjustable robust model by reformulating the homogeneity constraints according to the equivalence
@@
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (3a)
subject to $\min_{\boldsymbol{\phi}\in\mathcal{U}}\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (3b)
$\max_{\boldsymbol{\phi}\in\mathcal{U}}\left\{\phi_{v}d_{v}(x)-\mu\phi_{u}d_{u}(x)\right\}\leq 0,$ $u,v\in T$ (3c)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (3d)
@@
Given a radiosensitivity vector $\widehat{\boldsymbol{\phi}}$ that is estimated based on a patient’s imaging data at the start of the treatment, the goal is to construct an uncertainty set that satisfies our three modeling requirements, described in Section 2.2. Perhaps the simplest uncertainty set that can be applied is a box (in particular, a hypercube) uncertainty set, which allows the radiosensitivity of each PTV voxel $v$ to deviate from its calculated value $\widehat{\phi}_{v}$ by at most $\delta>0$, where $\delta$ is a given parameter. That is, $\mathcal{U}_{\text{B}}=\{\boldsymbol{\phi}\in\mathbb{R}^{|T|}:|\phi_{v}-\widehat{\phi}_{v}|\leq\delta,\;\forall v\in T\}$. This uncertainty set, used for example in Nohadani and Roy, (2017), admits a straightforward robust formulation. However, if $\delta$ is chosen to be sufficiently large to capture the actual uncertainty in the tumor biological conditions, then the resulting set may be overly conservative, thereby violating one of our key modeling requirements (see Section 2.2). To restrict the uncertainty further, a key observation to be made is that voxels that are situated close to each other should have similar radiosensitivities. In Figure 1, both plots demonstrate that the radiosensitivity differences are small for neighboring voxels and increase with the distance.
@@
Figure 1: Relationship between radiosensitivity difference and pairwise PTV voxel distance in the first-visit data of Patient 1 from the TCIA brain dataset (Clark et al., 2013).
Incorporating this observation into the uncertainty model leads to a spatially rbound (SB) uncertainty set; in particular, we define
@@
For any $u,v\in T$, $\gamma_{uv}=\Gamma(\operatorname{{dist}}(u,v))$, where $\operatorname{{dist}}(u,v)$ is the distance in voxels between $u$ and $v$, and $\Gamma:\mathbb{R}\rightarrow[0,1]$ is a nondecreasing subadditive function satisfying $\Gamma(x)=0$ if and only if $x=0$.
@@
The $\gamma_{uv}$ parameters for $u,v\in T$ can be personalized to a particular patient’s data, thereby addressing requirement (i). Requirement (iii) may be satisfied by imposing the spatial constraints associated with the parameters $\gamma_{uv}$, for $u,v\in T$, to augment the box constraints in the definition of the spatially bound uncertainty set. In the absence of these spatial constraints, the uncertainty set would reduce to the overly conservative box uncertainty set. The following section addresses the question of whether optimal solutions of (3) for uncertainty set $\mathcal{U}_{\text{SB}}$ can be computed efficiently in practice (requirement (ii)).